If I invest $5000 at 5% per year, how much money should I invest at 8% per year so that the annual income from the two investments is 6%?
Answer by richwmiller(17219)

0.08*x+0.05*5000=0.06(5000+x)
0.08*x+250=300+0.06x
0.08x-0.06x=300-250
0.02x=50
x=2500 at 8% to make 6% on 7500
check
0.08*2500+0.05*5000=0.06(5000+2500)
200+250=0.06(7500)
450=450
ok